OpenStudy (anonymous):

A function v is defined by @ganeshie8

OpenStudy (anonymous):

\[v:x \rightarrow \frac{ x-1 }{ x+1 },x \neq-1\]

OpenStudy (anonymous):

a)Find in similar form

OpenStudy (anonymous):

\[i)~v^{18}\]

ganeshie8 (ganeshie8):

isn't it just \(\left(\dfrac{x-1}{x+1}\right)^{18}\) ? are you given answer ?

myininaya (myininaya):

does similar mean in the form \[\frac{ax-b}{ax+c}\]

OpenStudy (anonymous):

That means that @ganeshie8 is right\[(\frac{ x-1 }{ x+1 })^{18}\] Thnx @ganeshie8 @myininaya

myininaya (myininaya):

oh ok

myininaya (myininaya):

I don't know need a medal I didn't do anything I just asked a question

myininaya (myininaya):

I wasn't saying what @ganeshie8 had was right I didn't know if there was some magical meaning behind similar or not :p

myininaya (myininaya):

but yeah normally v=(x-1)/(x+1) implies v^n=[(x-1)/(x+1)]^n I don't know what else it could mean

ganeshie8 (ganeshie8):

I think by similar form, they want you expand numerator and denominator

myininaya (myininaya):

eww

myininaya (myininaya):

you could divide the inside but you will end up with (18+1) terms

ganeshie8 (ganeshie8):

\[v^{18} : \left(\dfrac{x-1}{x+1}\right)^{18} = \dfrac{(1-x)^{18}}{(1+x)^{18}} = \text{you dont want to expand :/}\]

myininaya (myininaya):

this question seems weird i don't know find v^(18) given v similar form? it all sounds odd

myininaya (myininaya):

like I feel like I'm crazy for thinking that

myininaya (myininaya):

I seem some of his previous questions and they look harder than this

myininaya (myininaya):

seen*

ganeshie8 (ganeshie8):

15th question

myininaya (myininaya):

wait not to weird but do you think for example in this case v^2 means v composed with v

ganeshie8 (ganeshie8):

yeah it seems the exponent is refering to the number of times we need to compose..

myininaya (myininaya):

that would mean v^2=-1/x

ganeshie8 (ganeshie8):

Wow! thats it i guess @MARC_

myininaya (myininaya):

like @marc did have the answer as -1/x and he deleted it like I'm not saying that is the answer but that definitely looks more similar to the given form we have above

ganeshie8 (ganeshie8):

v^18 = v(v(v(v(... 18 times )))) ?

myininaya (myininaya):

lol I think so but I think we can try to find a pattern maybe for smaller numbers and make a guess for v^(18) based on that pattern

myininaya (myininaya):

\[v^1=\frac{x-1}{x+1} \\ v^2=\frac{\frac{x-1}{x+1}-1}{\frac{x-1}{x+1}+1}=\frac{x-1-(x+1)}{x-1+(x+1)}=\frac{-2}{2x}=\frac{-1}{x} \\ v^3=v(v(v(x))=v(\frac{-1}{x})=?\] @marc if you are still here do you want to try to do v(-1/x)?

myininaya (myininaya):

I think it actually starts over again at a certain number

OpenStudy (anonymous):

sorry for the late reply,\[=\frac{ -\frac{ 1 }{ x }-1 }{ -\frac{ 1 }{ x }+1 }\]

OpenStudy (anonymous):

\[=\frac{ \frac{ -1-x }{ x } }{ \frac{ x-1 }{ x } }\]

OpenStudy (anonymous):

\[=\frac{ -1-x }{ x }\div \frac{ x-1 }{ x }\]

OpenStudy (anonymous):

\[=\frac{ -1-x }{ x-1 }\]

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

yes that is v^3

myininaya (myininaya):

now to find v^4 plug v^3 result into the v function given

myininaya (myininaya):

\[v^4=v(v^3)=v(\frac{-1-x}{x-1})=?\]

myininaya (myininaya):

believe it or not this is going to look pretty very soon

OpenStudy (anonymous):

\[v^4=x\] @myininaya

myininaya (myininaya):

then you see you start over with v^5=v^1 and so therefore v^6=v^2 and v^7=v^3 and v^8=v^4 and v^9=v^5=v^1 and so on... do you see the pattern?

OpenStudy (anonymous):

i got it @myininaya Assume that every \[v^4\]is \[x\]

OpenStudy (anonymous):

\[v^4=x\]\[v^8=x\]\[v^{12}=x\]\[v^{16}=x\]\[v^{17}=\frac{ x-1 }{ x+1 }\]\[v^{18}=-\frac{ 1 }{ x }\]

OpenStudy (anonymous):

Thnx @myininaya